Solve the equation: x-3-sqrt of x-1=0
The answer is one of the following: A.x=1,7 B.x=2,5 C.x=5 D.x=7
Thank you for your time!

Answer by stanbon(75887)

Solve the equation: x-3-sqrt of x-1=0
sqrt(x-1)=x-3
Square both sides to get:
x-1=x^2-6x+9
x^2-7x+10=0
(x-5)(x-2)=0
x=5 or x=2
Checking these answers in the original equation
you will find that x=5 is a solution but x=2
is extraneous solution.
Cheers,
stan H.